# leftmargin = 20pt, rightmargin = 20pt, backgroundcolor = bggray, middlelinecolor = black, roundcorner = 15

### v1

@@ 5.1 Distributed Computing on the NOSTR Protocol
Service providers communicate with the customer through Job Result events, by providing the (optionally) encrypted output of the job. In the case of AI training, the output corresponds to an optimized version of the model parameters. The format of the Job Result event appears in the Event Type 3 above. The tag of the Job Result event include the original job request event id (stringified-JSON), the public key of the customer, amount that the Service Provider is requesting to be paid or an lightning invoice (bolt11 [25, 26]), and the optimized model parameter as an encrypted output. Finally, a tag "i" may be considered for additional information related to the initial job request or for validation of the output by the customer. {mdframed}
@@
During the model training process, service providers can communicate and provide feedback for an ongoing job through the Job Feedback event. The Job Feedback events may be considered in order to avoid time-outs and ensure that a service provider has achieved some partial progress. Information about partial progress can be optionally included by introducing additional fields in the tags of the event format in the Event Type 4. {mdframed}
The predefined format of the Job Feedback event includes the following tags [24]: {outline} \1 <content>: It may be empty, a final job-result or a partial-result (for a job in progress). This field will have an encrypted payload with p tag as key. \1 <amount>: Requested payment amount. \1 <status>: The service provider publishes the current status of the requested job. For instance: \2 payment-required: Service Provider requires payment before continuing. \2 processing: Service Provider is processing the job. \2 error: Service Provider was unable to process the job. \2 success: Service Provider successfully processed the job. \2 partial: Service Provider partially processed the job. The content filed might include a sample of the partial results.
@@
Service Providers have the option to announce availability or to advertise their support for specific job kinds by publishing discoverability events [27, (NIP-89)]. The format structure of the discoverability event appears in the Event Type 5. By including a field tag "i", service providers may provide additional information related to their available computational capabilities, hardware specification or execution time limits. {mdframed}
@@ 6 Payments
The NOSTR protocol has an integrated protocol flow for payments; requests for payment, invoice notifications, receipts for payments, as well as payments to certain events or directly to a user or a group of users. The payments take place through the lightning network [7] . For any form of application including social media, Internet-of-Things applications [28] and marketplaces [29], the lightning network enables bitcoin payments with instant settlement and virtually zero transaction fee. These properties together with in-channel messages are ideal for reliable handshaking mechanisms, anti-spam and on the fly payments between multiple job rounds among devices.
To further explain how the payments blend in with the rest of the protocol we proceed with an example. In the case of federated learning the aggregator itself may act as a customer who pays service providers to optimize a model on different subsets of a training data-set through multiple job rounds. Initially, the customer (aggregator) identifies possible service providers through the discoverability event (kind 31990, event type 5), selects a number of them and submits a Job Request event (kind , event type 2). The service providers respond with a Job Feedback event (kind , event type 4) with status payment-required, to announce that they are available to complete the job. Then the customer proceeds by submitting a small amount (of the total) as a payment and the service provider starts the job. Upon delivery of the final or partial output result (Job Result 3 & Job Feedback 4), the customer proceeds with a partial or total payment. This process repeats for each optimization round and for all service providers (assuming that they provide valid results). The ability to break down the payments into partial amounts and pay upon delivery of a valid result significantly minimizes the chance of stealing funds without completing the requested task. This strategy together with a reputation system can make the marketplace robust against adversarial or spam attacks up to a large extend. However, an in depth implementation of such sophisticated strategies remains out of the scope of this work. We discuss strategies for validation of the delivered output in Section 7.1.
@@ 6.1 Payments Implementation
Herein, we discuss the payment system which is integrated into the NOSTR protocol. We indicatively explain three components of the payment system; Payment Request, Payment Receipts and Validation of Receipts. For detailed documentation of the payment protocol flow, we refer the reader to implementation 57 [30, 31].
@@
A payment (known as zap) request is an event of kind 9734 that is not published to relays, but is instead sent to a lnurl pay callback url [32] of the recipient. The field content may be an optional message to send along with the payment. The event must include the following tags:
Relays is a list of relays for the wallet of the recipient to publish its payment (zap) receipt to.
Amount is the amount in millisats the sender intends to pay, formatted as a string.
Lnurl is the lnurl pay url of the recipient, encoded using bech32 [25] with the prefix lnurl.
p is the hex-encoded pubkey of the recipient.
e is an optional hex-encoded event id. Implementations must include this while paying an event rather than a public key.
@@
A payment receipt is created by a lightning node when an invoice generated by a payment request is paid. Payment receipts are only created when the invoice description (committed to the description hash) contains a payment request note. When receiving a payment, a series of steps are executed and a NOSTR event of kind 9735 (as appears below; Event Type 7) should be published to the relays declared in the payment request. We refer the reader to the detailed documentation [30] for a complete description of the Payment Receipt event structure and the full protocol flow.
@@
The service provider can retrieve payment receipts using a NIP- filter [33] and validate each receipt to verify that a customer fetched their invoice and claims payment submission for a certain job. Payments are validated by using the following steps:
The public key in the payment receipt event must be the same as the public key of the service provider.
The invoice Amount contained in the bolt11 tag of the payment receipt must be equal to the amount tag of the payment request.
The lnurl tag of the payment request should be equal the lnurl of the recipient.
We highlight the fact that a receipt is not a proof of payment, and it only proves that some NOSTR user fetched an invoice. As a final step of payment validation, the recipient must verify the payment on their lightning node or payment processor.
@@ 7 Money-In AI-Out: Marketplace for Federated Learning and LLMs
Herein, we introduce the protocol flow and algorithmic design for federated learning and LLMs training over the NOSTR protocol. By leveraging the existing structure of the protocol for data processing, and by introducing additional components when it is necessary, we construct an open and decentralized marketplace for training AI models. In this market, a user (customer) provides a dataset, model specifications, the service providers (AI Vending Machines) receive a payment, and they return an AI model trained on this dataset. Customers and service providers communicate through relays, which can be public or private (and self hosted). Also, the communication can be implemented with encrypted information to guarantee privacy for the dataset or for the output of the computation. We proceed by explaining further the protocol flow and then we present a detailed algorithmic implementation for customers and service providers.
@@
We propose a protocol flow that involves multiple job rounds between the customer and service providers. Through multiple job rounds we ensure that there is a valid progress in the optimization process by each of the service providers individually. This allows to consider on the fly payments after validation of the output at each round. If the service providers give an invalid output or if they delay the job beyond a certain amount of time, then the customer proceeds to reassign the job to another service provider without paying for the invalid (or delayed) job output.
@@
In this work we consider two training methods: the classical federated learning [34, FEDAVG] and DiLoCo algorithm [8] for training LLMs. Both of these have an inner and outer optimization part. This allows us to present our approach in parallel. Specifically, we define the job round as the sequence of the following successful publication of events and actions; for the first part (inner optimization): job request, inner optimization computation executed by the service providers, retrieval of the outputs from inner optimization routines by the customer through job result events, validation of the output result. For the second part (Outer optimization): use the latest inner optimization outputs as inputs and execute the outer optimization (through a job request, if needed or locally directly executed by the customer), complete any pending payments and update the model parameters for the next round (if any). Upon validation of the outputs a (partial) payment is submitted by the customer to each of the service providers. To explain this further we provide the sequence of the steps below.
Protocol Flow.
At every job round: {outline} \1 Customer (Algorithm 1) publishes a job request (e.g. kind:8000, Federated Learning, inner optimization). The job request may include a list of chosen service providers, and it must include all the required information: the current state of the model (for instance for the first round initial value of model parameters), a link (source) to the data, list of relays where the service providers should publish their job results job feedback at, as well as all the required parameters (see Event Type 2). \1 Service Providers (Algorithm 4) must submit job-feedback events (kind:7000, e.g. partial payment-required) before they start the job with STATUS = payment_required, if they wish a (partial) payment before starting computation (Event Type 4). This requests a partial payment and signals that a service provider is available to start upon receiving the payment. \1 Customer submits a initial payment to service providers (responsible for inner optimization). \1 Upon receiving the partial payment the service providers start executing the inner optimization job (Algorithm 5). \2 Within a fixed time period service providers publish Job Feedback events (Event Type 4) with partial/sample result. \2 If results delivery exceeds a predefined time limit, then customer stops the procedure and seeks for other service providers (responsible for inner optimization). \1 Upon completion, each service provider publishes the output of the job through a Job Result event (kind:6000) (Event Type 3). \1 Customer validates the results/ouputs (Algorithm 3) \2 If the result/ouput is valid then the customer completes the payments to all service providers (responsible for inner optimization) with valid output. \2 If the result is not valid then reassigns the job to a different service provider (Algorithm 2). \1 With successful outputs for all inner optimization jobs: The customer proceeds to the outer optimization (Algorithm 6). \2 If the customer selects to execute the outer optimization locally, then the customer executes the outer optimization (Algorithm 6) with input the output of inner optimization procedures. \2 Else if the customer selects to assign the task to a service provider for the outer optimization, then the customer follows the steps of reassigning the job from SELF to a service provider (Algorithm 2), retrieving and validating the output, and completing any payments identically to the inner optimization job handling as was described above.
At the end of each round the customer updates the model parameters to use them as input for the next round. \2 If the current round is the terminal round then the customer completes any pending payments and ends the procedure with output the final result. \1 At any point, if there is an amount pending to be paid as instructed by the service provider, the customer can pay the lightning-invoice [25, 26, bolt11] (see lnurl Event Type 6). We present an implementation of the protocol flow in Section 7.1 and Section 7.2.
@@ 7.1 Customer
A complete implementation of the customer routine appears in Algorithm 1. This includes the steps that the customer follows and the events that the customer publishes. First, the customer discovers and requests a set of service providers to run inner optimization part of the job to as we explained in the protocol flow above. Upon successful delivery of outputs, the customer proceeds to the outer optimization. The outer optimization for both FedAvg and DiLoCo is a single or double update of local variables respectively. However we consider the option for the customer to also assign this computation step to a service provider. This can be useful in practice if someone prefers to implement an outer optimizer with higher computational complexity (for instance multiple iterations) and the customer is not able to perform such computations locally. We proceed by explaining further three parts of the main customer routine: how the customer handles disconnections and failures of a service provider to respond (Algorithm 1), how the customer validates the output results given by a service provider at job completion (Algorithm 3), and how the customer reassigns a job to a different service provider if needed (Algorithm 2).
@@
The customer may consider a time-out period. If a service provider did not provide the output within a certain time window (from the time point of receiving the initialization payment), then the customer proceeds to reassign the job to a different service provider (Algorithm 2). As an example of the time-out implementation see Algorithm 1 (lines 13-14). All the information regarding the time-out rules should be included in the Job Request event (Event Type 2).
The definition of failure to respond may vary from one implementation to another. As an example, the customer may choose to receive a valid output within a certain time period. Alternatively the time-out decision may depend on the periodic responses by the service provider through Job Feedback events (Event Type 4), that notify the customer about the status of the job and provide partial results of the computation.
@@
Upon receiving a Job Result event (Event Type 3) the customer proceeds to verify the accuracy of the output. In the context of training AI models with FEDAVG and DiLoCl, we consider two conditions that test the decay of the loss on a validation dataset. Specifically, Algorithm 3 requires as input the validation dataset, the current state of the model (model parameters), and a threshold for the two following policies. As a first option, we compare the loss decay progress relative to other service providers (see accuracy against service providers, Algorithm 3). This validation test has been introduced in [35, Accuracy Checking]. As a second option, we consider a test that checks if a moving average of loss decays with a certain rate (Algorithm 3). Alternative accuracy tests and approaches to identify malicious service providers appear in prior works [36, 35, 37, 38]. Such implementations may be considered in parallel with our design but remain out of the scope of this work.
@@
Under a time-out or a failed validation instance, the customer reassigns the job to a different service provider. We provide an example of such implementation in Algorithm 2. Further, under perpetual time-outs and invalid output detection, the customer continues to reassign the job, until a service provider delivers a valid output. Then the routine returns the public key of new service provider and the customer updates the list of the service providers under consideration.

### v2

@@ 5.1 Distributed Computing on the NOSTR Protocol
Service providers communicate with the customer through Job Result events, providing the (optionally encrypted) output of the job. In the case of AI training, the output corresponds to an optimized version of the model parameters. The format of the Job Result event appears in the Event Type 3 above. The tags of the Job Result event include the original job request event id (stringified-JSON), the public key of the customer, amount that the Service Provider is requesting to be paid or an lightning invoice (bolt11 [25, 26]), and the optimized model parameter as an encrypted output. Finally, an "i" tag provides additional information related to the initial job request or for validation of the output by the customer.
"pubkey": "<service-procider pubkey>",
"content": "<enctypted payload>",
"kind": 6xxx // kind in range 6000 - 6999,
"tags":
Event Type 3: Job Result Event
@@
During the model training process, service providers can communicate feedback for an ongoing job through the Job Feedback event. Customers may demand job feedback to avoid timeouts and ensure that a service provider has achieved some partial progress. Information about partial progress can be optionally included by adding fields to the tags in Event Type 4.
"kind": 7000,
"content": "<empty-or-payload>",
"tags":
Event Type 4: Job Feedback Event
The predefined format of the Job Feedback event includes the following tags [24]:
<content>: It may be empty, a final job result, or a partial result (for a job in progress). This field will have an encrypted payload with p tag as key.
<amount>: Requested payment amount.
<status>: The service provider publishes the current status of the requested job. For instance:
payment-required: Service Provider requires payment before continuing
processing: Service Provider is processing the job
error: Service Provider was unable to process the job
success: Service Provider successfully processed the job
partial: Service Provider partially processed the job. The content filed might include a sample of the partial results.
@@
Service Providers can announce their availability or to advertise their support for specific job kinds by publishing discoverability events [27, (NIP-89)]. The format structure of the discoverability event appears in the Event Type 5. By including an "i" tag, service providers may provide additional information related to their available computational capabilities, hardware specifications, or execution time limits.
"kind": 31990,
"pubkey": "<pubkey>",
"content": [
\"name\": "Federated Learning AI-VM" ,
\"about\": "I’m a AI-VM for federated learning."
"tags":
Event Type 5: Discoverability Event
@@ 6 Payments
The NOSTR protocol includes integrated payment functionality; requests for payment, invoice notifications, receipts for payments, as well as payments to certain events or directly to a user or a group of users. The payments take place through the lightning network [7]
. For any form of application, including social media, Internet-of-Things applications [28], and marketplaces [29], the lightning network enables bitcoin payments with instant settlement and virtually zero transaction fees. These properties, together with in-channel messages, are ideal for reliable handshaking mechanisms, anti-spam and on the fly payments between multiple job rounds among devices.
To further explain how the payments blend in with the rest of the protocol, we proceed with an example. In the case of federated learning, the aggregator itself may act as a customer who pays service providers to optimize a model on different subsets of a training dataset through multiple job rounds. Initially, the customer (aggregator) identifies possible service providers through the discoverability event (kind 31990, Event Type 5), selects a number of them and submits a Job Request event (kind , event type 2). The service providers respond with a Job Feedback event (kind , event type 4) with status payment-required, to announce that they are available to complete the job. Then the customer proceeds by submitting a small amount (of the total) as a payment and the service provider starts the job. Upon delivery of the final or partial output result (Job Result 3 & Job Feedback 4 events), the customer proceeds with a partial or total payment. This process repeats for each optimization round and for all service providers (assuming that they provide valid results). The ability to break down payments into partial amounts and pay upon delivery of a valid result significantly reduces the possibility that payments will be executed without completing the requested task. This strategy, together with a reputation system, can make the marketplace robust against adversarial or spam attacks to a large extent. However, an in-depth implementation of such sophisticated strategies remains out of the scope of this work. We discuss strategies for validation of the delivered output in Section 7.1.
@@ 6.1 Payments Implementation
In this section, we discuss the payment system which is integrated into the NOSTR protocol. We briefly explain three components of the payment system; Payment Request, Payment Receipts and Validation of Receipts. For detailed documentation of the payment protocol flow, we refer the reader to implementation 57 [30, 31].
@@
A payment (known as zap) request is an event of kind 9734 that is not published to relays, but is instead sent to a lnurl pay callback url [32] of the recipient. The field content may be an optional message to send along with the payment. The payment request event must include the following tags:
Relays: A list of relays where the recipient’s wallet should publish its zap receipt.
Amount: The payment amount in millisats that the sender intends to pay, formatted as a string.
Lnurl: The lnurl pay url of the recipient, encoded using bech32 [25] with the prefix lnurl.
p: The hex-encoded pubkey of the recipient.
e: An optional hex-encoded event ID. Implementations must include this when paying for an event rather than directly to a public key.
"kind": 9734,
"content": "Zap!",
"tags":
"relays", "<wss://relay-domain>", "wss://anotherrelay.com",
"amount", "<msats-amount>",
"lnurl", "<a-static-lightning-invoice>",
"p", "<hex-encoded-pubkey-of-the-recipient>",
"e", "<optional-hex-encoded-event-ID>"
"pubkey": "<NPUB>",
"created_at": <timestamp>,
"id": "<event-ID>",
"sig": "<signature>"
Event Type 6: Payment Request Event
@@
A payment receipt is created by a lightning node when an invoice generated by a payment request is paid. Payment receipts are only created when the invoice description (which is committed to the description hash) contains a payment request note. When receiving a payment, a series of steps is executed, and a NOSTR event of kind 9735 (shown below as Event Type 7) should be published to the relays declared in the payment request. We refer the reader to the detailed documentation [30] for a complete description of the Payment Receipt event structure and the full protocol flow.
"id": "<event-ID>",
"pubkey": "<NPUB>",
"created_at": <paid-at-date>,
"kind": 9735,
"tags":
"p", "<payment-recipient)>",
"P", "<optional-P-tag-from-the-pubkey-of-the-payment-request>",
"e", "<optional-tag-same-as-payment-request-e-tag>",
"bolt11", "<lightning-invoice>",
"description", "<SHA256(description)-from-invoice>",
"preimage", "<payment-preimage>"
"content": "",
Event Type 7: Payment Receipt Event
@@
The service provider can retrieve payment receipts using a NIP- filter [33] and validate each receipt to verify that a customer fetched their invoice and their payment claim for a certain job. Payment validation follows these steps:
The public key in the payment receipt event must match the public key of the service provider.
The invoice amount contained in the bolt11 tag of the payment receipt must equal the amount in the amount tag of the payment request.
The lnurl tag of the payment request should be equal to the lnurl of the recipient.
We highlight that a receipt is not proof of payment; it only proves that some NOSTR user fetched an invoice. As a final validation step, the recipient must verify the payment on their Lightning node or payment processor.
@@ 7 Money-In AI-Out: Marketplace for Federated Learning and LLMs
Herein, we introduce the protocol flow and algorithmic design for federated learning and LLMs training over the NOSTR protocol. By leveraging the existing structure of the protocol for data processing, and by introducing additional components when it is necessary, we construct an open and decentralized marketplace for training AI models. In this market, a customer provides a dataset and model specifications. The service providers (AI Vending Machines) receive a payment, and return an AI model trained on this dataset. Customers and service providers communicate through relays, which can be public or private (and self hosted). Additionally, the communication can use encryption to guarantee privacy for the dataset or the computational output. We proceed by explaining further the protocol flow and then we present a detailed algorithmic implementation for customers and service providers.
@@
We propose a protocol flow involving multiple job rounds between customers and service providers. Through multiple job rounds, we ensure valid progress in the optimization process of each service provider individually. This enables on-the-fly payments after validation of the output at each round. If a service provider delivers an invalid output or delays the job beyond a certain time limit, the customer reassigns the job to another service provider without paying the invalid (or delayed) job output.
@@
In this work, we consider two training methods: the classical federated learning [34, FEDAVG] and the DiLoCo algorithm [8] for training LLMs. Both of these have an inner and outer optimization phases. This allows us to present our approach in parallel. Specifically, we define the job round as the successful sequence of the following events and actions. For the first part (inner optimization): job request, inner optimization computation executed by the service providers, retrieval of outputs by the customer through job result events, validation of the outputs. For the second part (outer optimization): use the latest inner optimization outputs as inputs, execute the outer optimization (either through a job request, or executed locally by the customer), complete any pending payments and update the model parameters for the next round (if any). Upon validation of the outputs, a (partial) payment is submitted by the customer to each service provider. To explain this further we provide the sequence of the steps below.
Protocol Flow for Each Job Round.
At every job round:
Customer (Algorithm 1) publishes a job request (e.g. kind 8000: Federated Learning, inner optimization). The job request may include a list of chosen service providers, and it must include all the required information: the current state of the model (for instance for the first round initial value of model parameters), a link (source) to the data, a list of relays where the service providers should publish their job results job feedback at, as well as all the required parameters (see Event Type 2).
Service Providers (Algorithm 4) must submit job feedback events (kind 7000, e.g. partial payment required) before they start the job with STATUS = payment_required, if they wish a (partial) payment before starting computation (Event Type 4). This requests a partial payment and signals that the service provider is ready to start upon receiving it.
Customer submits an initial payment to service providers (responsible for inner optimization).
Upon receiving the partial payment the service providers start executing the inner optimization job (Algorithm 5).
Within a fixed time period, service providers publish Job Feedback events (Event Type 4) with partial/sample results.
If results delivery exceeds a predefined time limit, the customer stops the procedure and seeks other service providers (responsible for inner optimization).
Upon completion, each service provider publishes the output of the job through a Job Result event (kind 6000, Event Type 3).
Customer validates the results/ouputs (Algorithm 3)
If the output is valid, then the customer completes payments to all service providers (responsible for inner optimization) who returned valid outputs.
If the result is not valid, the customer reassigns the job to a different service provider (Algorithm 2).
Upon receiving successful outputs for all inner optimization jobs, the customer proceeds to the outer optimization (Algorithm 6).
If the customer selects to execute the outer optimization locally, then the customer executes the outer optimization (Algorithm 6) with input the output of inner optimization procedures.
Otherwise, if the customer chooses to assign the task to a service provider for the outer optimization, the customer follows the steps for reassigning the job from SELF to a service provider (Algorithm 2), retrieving and validating the output, and completing any payments identically to the inner optimization job handling as described above.
At the end of each round, the customer updates the model parameters to use them as input for the next round.
If the current round is the terminal round, then the customer completes any pending payments and ends the procedure outputting the final result.
At any point, if there is an amount pending payment as instructed by the service provider, the customer can pay the Lightning-invoice [25, 26, bolt11] (see lnurl Event Type 6).
We present an implementation of the protocol flow in Section 7.1 and Section 7.2.
@@ 7.1 Customer
A complete implementation of the customer routine appears in Algorithm 1. This includes the steps followed and the events published by the customer. First, the customer discovers and requests a set of service providers to run the inner optimization part of the job, as explained in the protocol flow above. Upon successful delivery of outputs, the customer proceeds to the outer optimization. The outer optimization for both FedAvg and DiLoCo is a single or double update of local variables respectively. However we consider the option for the customer to also assign this computation step to a service provider. This can be useful in practice if a customer prefers to implement an outer optimizer with higher computational complexity (for instance, multiple iterations) but is unable to perform such computations locally. We now explain three parts of the main customer routine in detail: (1) how the customer handles disconnections and failures to respond (Algorithm 1), (2) how the customer validates the output results provided by a service provider at job completion (Algorithm 3), and (3) how the customer reassigns a job to a different service provider if needed (Algorithm 2).
@@
The customer may consider a timeout period. If a service provider does not provide the output within a certain time window (from receiving the initial payment), the customer reassigns the job to a different service provider (Algorithm 2). As an example of the timeout implementation, see Algorithm 1 (lines 13-14). All information regarding the timeout rules should be included in the Job Request event (Event Type 2).
The definition of failure to respond may vary from one implementation to another. As an example, the customer may choose to receive a valid output within a certain time period. Alternatively, the timeout decision may depend on periodic responses from the service provider through Job Feedback events (Event Type 4), which notify the customer about the status of the job and provide partial computation results.
@@
Upon receiving a Job Result event (Event Type 3), the customer proceeds to verify the accuracy of the output. In the context of training AI models with FEDAVG and DiLoCo, we consider two conditions that test the loss decay on a validation dataset. Specifically, Algorithm 3 requires as input the validation dataset, the current state of the model (model parameters), and a threshold for the two following policies. As a first option, we compare the loss decay relative to other service providers (see Accuracy Against Service Providers, Algorithm 3). This validation test has been introduced in [35, Accuracy Checking]. As a second option, we consider a test that checks whether a moving average of the loss decays at a certain rate (Algorithm 3). Alternative accuracy tests and approaches for identifying malicious service providers appear in prior works [36, 35, 37, 38]. Such implementations could be integrated with our design but remain outside the scope of this work.
Algorithm 1 Customer
1:C_NPUB: Customer’s NOSTR public key, DATA_SOURCE: link to the data, NUM_PR: # of providers, NUM_JOBS : Total # of jobs for each provider, RELAYS: a set of relays
2:Connection through websockets to all relays in RELAYS Where events are published at
3:Search Routine: Search for NUM_PR AVAILABLE service providers. Event Type 5
4:Search Routine returns: PROVIDERS with NUM_PR NPUBs entries
5:Split the data into NUM_PR parts and generate the list DATA_SOURCES with NUM_PR sources for each data segment DATA_SOURCES is the data source of the provider
6:Initialize the optimization parameters and model: model_parameters, model, …
7:Publish Job Requests KIND: 8000 with all the required information Event Type 2
8:Wait for Job Feedback KIND: 7000 and payment-request from service providers Event Type 4
9:Submit an initialization payment to PROVIDERS
10:for job in (1,NUM_JOBS) do
11: Initialize done FALSE
12: while done TRUE do
13: Wait some time
14: Fetch Job Feedback events (KIND:7000) from each provider in PROVIDERS
15: if timeout occurs then
16: Call Reassign (Algorithm 2) to replace all the delayed service providers, update PROVIDERS
17: if JOB_STATUS = success for all providers in PROVIDERS then
done = TRUE
18: Fetch Job Results (KIND: 6000) and "outputs" from all providers in PROVIDERS Event Type 3
19: for service_provider in PROVIDERS call the Validation routine (Algorithm 3) and do
20: while Validation_Test = FAIL for the service_provider do
21: Reassign the job of service_provider; call Reassign (Algorithm 2), update PROVIDERS
22: Submit (partial) payments to service_provider and publish Payment Receipt (KIND: 9735)
23: if OUTER_OPTIMIZER SELF then
24: Run the Outer_Optimization routine (Algorithm 6) with input the current updates "outputs"
25: else
26: Reassign the Outer Optimization from SELF to a service provider (Algorithm 2)
27: Submit (partial) payment to OUTER_OPTIMIZER and publish Payment Receipt (KIND: 9735)
28: Update model_parameters using the "output" of the OUTER_OPTIMIZER
29: if stopping condition is satisfied then
30: break
31: else if job NUM_JOBS then
32: Publish Job Requests KIND: 8000 with all the required information Event Type 2
33:Finalize payments to every service_provider in PROVIDERS and publish receipts Event Type 7
34:return model_parameters
@@
Under a timeout or a failed validation instance, the customer reassigns the job to a different service provider. We provide an example of such implementation in Algorithm 2. Further, under perpetual timeouts and invalid output detection, the customer continues to reassign the job, until a service provider delivers a valid output. Then the routine returns the public key of new service provider and the customer updates the list of the service providers under consideration.
Algorithm 2 Reassign the Job of a Service Provider to a Different Service Provider
1:SP_NPUB: public key of the service provider that will be replaced, <computation information & time>, RUN_OPTION: <FEDAVG or DiLoCo>, DATA: DATA_SOURCES , PROVIDERS, TASK: <INNER or OUTER>
2:Connection through websockets to all relays in RELAYS Where events are published at
3:Search Routine: Search for an AVAILABLE service provider for KIND: 8000. Event Type 5
4:Search Routine returns: A single public key: SP_NPUB_NEW
5:Publish a Job Request KIND: 8000: identical to the last job request for the service provider SP_NPUB, but with updated field "p" set to SP_NPUB_NEW Event Type 2
6:Wait for Job Feedback KIND: 7000 and payment-request from service provider SP_NPUB_NEW
7:Submit an initialization payment to SP_NPUB_NEW
8:while JOB_STATUS success for SP_NPUB_NEW do
9: Wait some time
10: Fetch Job Feedback events (KIND:7000) from service provider SP_NPUB_NEW Event Type 4
11: if timeout occurs then break and reassign again to replace SP_NPUB_NEW
12: if JOB_STATUS = success for SP_NPUB_NEW then
13: Fetch Job Result (KIND: 6000) and "output" from SP_NPUB_NEW Event Type 3
14: Call the Validation Routine
15: if Validation_Test = PASS then
16: return SP_NPUB_NEW, "output"
17: else if Validation_Test = FAIL then break and
reassign again to replace SP_NPUB_NEW
Algorithm 3 Validation of the Service Provider Output
1:SP_NPUB: service provider public key, RUN_OPTION: <FEDAVG or DiLoCo>, validation dataset: , model_parameters: , the last "outputs" of the Inner Optimization jobs: and the increments , policy thresholds: , TEST_TYPE
2:A history of model parameters is available. (TEST_TYPE = "B" requires the previous model parameters for the test condition)
3:if TEST_TYPE = "A" then Check accuracy against other service providers [35]
4:
5:
6: if then
7: SP_NPUB is malicious
8: return FAIL
9:else if TEST_TYPE = "B" then Check progress over time (with time lag)
10: , for all For Outer: replace with
11: if then Check if a moving average of the loss decays
12: return FAIL
13:return PASS